# Bridging Compositional and Distributional Semantics: A Survey on Latent Semantic Geometry via AutoEncoder

### v1

@@ 1 Introduction
In the field of natural language learning and generation, improving the geometric structure of semantic features within the representation space can enhance the controllability and interpretability of language models (LMs), which has been widely explored in the task, such as style-transfer Liu et al. (2023); Gu et al. (2023, 2022); Hu and Li (2021); Vasilakes et al. (2022). Those studies provide a quasi-symbolic, localised, compositional generation control to informal, less-rigid semantic features, such as sentiment and topic. Recently, a line of studies investigated this kind of control over formal semantic features Zhang et al. (2024a, 2025a, d, b, c, 2025b). Formal/symbolic semantics, such as Montague Semantics Dowty et al. (2012), Semantic Role Labelling Palmer et al. (2010), and Argument Structure Theory Jackendoff (1992), which provides a canonical, granular, and rigid representation of meaning in natural language, has been a central focus of philosophical and linguistic inquiry for thousands of years. Its goal is to systematically capture the logical structure of linguistic expressions and to model how complex meanings arise from simpler parts. All these studies provide valuable insights to bridge the compositional and distributional semantics.
@@
Composition and Generalisation: One of the enduring strengths of compositional semantics is its ability to handle compositionality—the principle that the meaning of a complex expression is determined by the meanings of its parts and the rules used to combine them. Embedding this principle into language models enhances their ability to generalise to novel combinations and linguistic constructions, beyond those seen during training. In the Math Reasoning task Meadows et al. (2024), for example, the model is tasked with predicting the output of a mathematical transformation given a specified operation and an input formula during training. However, current LMs often fail to generalise effectively to variable or length out-of-distribution (OOD) inputs Zhang et al. (2024d). Structured semantic representations—such as syntax trees that explicitly encode the function, constant, and variable—can provide a more robust framework for generalisation, enabling models to maintain compositional integrity even in the face of lexical variation.
@@
Syllogistic NLI Rule-based Explanatory NLI corpus Zhang et al. (2024c)
@@
IMDb movie review Maas et al. (2011)
Topic Style-transfer AGNews dataset Zhang et al. (2015)
Toxic Style-transfer Jigsaw Toxic Comment  [^1] https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge/
@@ 3 Data Resources
In addition to the two main types of expressions we focus on in compositional formal semantics, the NLP domain includes a wide range of corpora that explore compositional informal semantics across various downstream tasks, such as style transfer and paraphrasing. These datasets often capture nuanced variations in meaning associated with attributes such as sentiment, formality, or tone, and serve as valuable resources for investigating how language models handle semantic composition in more flexible and context-dependent settings. Typical corpora include Yelp reviews Shen et al. (2017), Yahoo answers Yang et al. (2017), dSentences  [^3] https://github.com/mcharrak/discreteVAE, IMDb movie review Maas et al. (2011), AGNews dataset Zhang et al. (2015), and Jigsaw Toxic Comment Classification Challenge Dataset  [^4] https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge/.
@@ 4.1.1 Transformer Language VAEs
Latent injection setup & Math formula Summarisation
KV Addition: softmax(Q [z + K]Td) [z + V] Simple and global interventions in the sequence of attention of LM decoder, which can mitigate the KL vanishing without additional parameters Zhang et al. (2024d).
@@
(2) KV Addition Injection. In addition to memory setup, $z$ can also be directly incorporated into the key and value matrices of the self-attention mechanism, as demonstrated in Zhang et al. (2024d).
@@
(3) Query Addition Injection. $z$ can also be directly added into the Query matrices of the self-attention mechanism for effectively search the pretrained, memorised knowledge from Key and Value by incorporating prior information Morris et al. (2025). For example, Zhang et al. (2024d) inject syntax information into Query for effective searching the semantic information from Key and Value.
@@ 4.1.2 Latent Space Geometry
(3) Syntax-Semantic Attribute Space. Several studies have explored methods for learning latent spaces in which syntactic structure and lexical semantics are disentangled. For example, Zhang et al. (2019) used two separate encoders to independently encode semantic and syntactic information. In contrast, Bao et al. (2019) utilised a single encoder within a VAE framework to disentangle semantic and syntactic representations through multi-task optimisation. Additionally, Zhang et al. (2024d) employed a graph-based encoder alongside a semantic encoder to explicitly capture syntax and semantics, mitigating the information bottleneck from a single encoder.
(4) Semantic Attribute Space. Defining generative factors in terms of semantic roles has recently been explored as a promising direction for structuring latent spaces. Carvalho et al. (2023) defined the generative factors of natural language definition, named Definition Semantic Roles (DSR) Silva et al. (2018), to assist the Definition Modelling tasks. Recent studies Zhang et al. (2024a, 2025a) formalise the latent space geometry from the perspective of Argument Structure Theory Jackendoff (1992); Levin (1993); Rappaport Hovav and Levin (2008). In this framework, each semantic role–word content combination is represented as a convex cone within the latent space, enabling localised semantic control and enhancing interpretability.
(5) Reasoning Attribute Space. More recent studies have also explored the encoding of reasoning patterns within the latent space of VAEs for Natural Language Inference (NLI) tasks Zhang et al. (2024c, 2025b). They seek to structure the latent space in a way that captures inferential relationships, such as mathematical reasoning rules, thereby enabling the model to represent not only the meanings of individual sentences but also the logical relations that hold between them.
@@ 4.2.1 Vector Quantized VAEs
(1) Topic Attribute Space. Topic-VQ-VAE Yoo and Choi (2024) focuses on document-level generation by learning topic-aware latent codes. It leverages vector quantization to discover discrete topic representations that guide the generation process, enabling more coherent and thematically consistent document outputs.
@@
(4) Semantic Space. To investigate the geometrical properties of discrete latent spaces, a recent study by Zhang et al. (2024b) integrates the VQ-VAE framework into the T5 encoder-decoder architecture to learn a discrete semantic space. The authors further propose techniques such as traversal and interpolation to explore and manipulate representations within this discrete space, targetting fine-grained semantic control.
@@ 5.2 Generation Evaluation
The interpolation mechanism operates by selecting two input sentences, $x_{1}$ and $x_{2}$, and encoding them into their corresponding latent representations $z_{1}$ and $z_{2}$. A linear interpolation is then performed along the latent space trajectory defined by $z_{t}=z_{1}\cdot(1-t)+z_{2}\cdot t$, where $t$ varies from 0 to 1 in increments of 0.1. Consequently, nine intermediate sentences are generated along the interpolation path. Prior work has shown that improved latent space geometry leads to smoother and more semantically coherent interpolations Zhang et al. (2024b).
@@
In VAE, for example, given two sentences that share the concept animal—animals require food for survival and animals require warmth for survival—their vector addition in the latent space may yield a new sentence such as animals produce milk, which also preserves the concept animal. This suggests the existence of a “convex cone” corresponding to the semantic concept animal within the latent space Zhang et al. (2024b, 2025a). Moreover, using arithmetic to steer the task-specific generative factor, such as sentiment, has been widely deployed in language VAE domain Shen et al. (2020); Mercatali and Freitas (2021). Furthermore, controlling (Large) LMs generation through the arithmetic of steering vectors to internal representations has been widely explored Turner et al. (2023), such as Concept Activation Addition (CAA) Rimsky et al. (2024), Representation Perturbation (RePe) Zou et al. (2023), Top PC Im and Li (2025), and Inference-Time Intervention (ITI) Li et al. (2023). In this context, the discrete concept vectors learned by SAEs can be interpreted as steering vectors, enabling targeted modulation of language model outputs Chalnev et al. (2024), such as SPARE Zhao et al. (2025) and SAE-SSV He et al. (2025).
@@ 6 Conclusion and Open Questions
Emerging diffusion-based LMs, such as Neuro-symbolic diffusion Nie et al. (2025); van Krieken et al. (2025), offer new opportunities to combine the generative strengths of large-scale models with symbolic-level controllability. However, the design of semantically interpretable latent priors within these frameworks is still in its early stages. Incorporating the compositional semantic representation can potentially improve the sampling efficiency during the denoising process. Developing structured latent representations that align with formal semantics could enable more transparent reasoning and controllable generation in future generative architectures.
@@ Appendix A Additional Examples
Figure 4: Generation evaluation for VAE or VQVAE latent space. More results can be found in Zhang et al. (2024a) for VAE and Zhang et al. (2024b) for VQVAE.

### v2

@@ 1 Introduction
In the field of natural language learning and generation, improving the geometric structure of semantic features within the representation space can enhance the controllability and interpretability of language models (LMs), which has been widely explored in the task, such as style-transfer Liu et al. (2023); Gu et al. (2023, 2022); Hu and Li (2021); Vasilakes et al. (2022). Those studies provide a quasi-symbolic, localised, compositional generation control to informal, less-rigid semantic features, such as sentiment and topic. Recently, a line of studies investigated this kind of control over formal semantic features Zhang et al. (2024b, 2025a, e, c, d, 2025b). Formal/symbolic semantics, such as Montague Semantics Dowty et al. (2012), Semantic Role Labelling Palmer et al. (2010), and Argument Structure Theory Jackendoff (1992), which provides a canonical, granular, and rigid representation of meaning in natural language, has been a central focus of philosophical and linguistic inquiry for thousands of years. Its goal is to systematically capture the logical structure of linguistic expressions and to model how complex meanings arise from simpler parts. All these studies provide valuable insights to bridge the compositional and distributional semantics.
@@
Composition and Generalisation: One of the enduring strengths of compositional semantics is its ability to handle compositionality—the principle that the meaning of a complex expression is determined by the meanings of its parts and the rules used to combine them. Embedding this principle into language models enhances their ability to generalise to novel combinations and linguistic constructions, beyond those seen during training. In the Math Reasoning task Meadows et al. (2024), for example, the model is tasked with predicting the output of a mathematical transformation given a specified operation and an input formula during training. However, current LMs often fail to generalise effectively to variable or length out-of-distribution (OOD) inputs Zhang et al. (2024e). Structured semantic representations—such as syntax trees that explicitly encode the function, constant, and variable—can provide a more robust framework for generalisation, enabling models to maintain compositional integrity even in the face of lexical variation.
@@
Syllogistic NLI Rule-based Explanatory NLI corpus Zhang et al. (2024d)
@@
IMDb movie review Diao et al. (2014); Maas et al. (2011); Li et al. (2018)
Stanford Sentiment Treebank Socher et al. (2013)
Formality Style-transfer GYAFC Rao and Tetreault (2018)
Codeswitching Style-transfer LinCE Aguilar et al. (2020)
Topic Style-transfer AGNews dataset Zhang et al. (2015), DBpedia Zhang et al. (2015)
Toxic Style-transfer Jigsaw Toxic Comment
@@ 3 Data Resources
In addition to the two main types of expressions we focus on in compositional formal semantics, the NLP domain includes a wide range of corpora that explore compositional informal semantics across various downstream tasks, such as style transfer and paraphrasing. These datasets often capture nuanced variations in meaning associated with attributes such as sentiment, formality, or tone, and serve as valuable resources for investigating how language models handle semantic composition in more flexible and context-dependent settings. Typical corpora include Yelp reviews Shen et al. (2017), Yahoo answers Yang et al. (2017), dSentences  [^2] https://github.com/mcharrak/discreteVAE, IMDb movie review Maas et al. (2011), AGNews dataset Zhang et al. (2015), and Jigsaw Toxic Comment Classification Challenge Dataset  [^3] https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge/. More style-transfer corpora can be found in the survey Jin et al. (2022).
{forest}
for tree= grow=east, draw, rounded corners, node options=align=center, font=, edge=->, >=latex, parent anchor=east, child anchor=west, l sep+=10pt, s sep+=5pt, anchor=base west, [Latent Semantic Geometry [Sparse AutoEncoder, fill=yellow!50 [Informal Linguistic Attribute Space, fill=yellow!20 [Elhage et al. (2022); Cunningham et al. (2023); Deng et al. (2025); He et al. (2025)] ] [Formal Linguistic Attribute Space, fill=yellow!20 [Jing et al. (2025)] ] ] [Vector Quantised VAE, fill=cyan!60 [Semantic or Concept Space, fill=cyan!30 [Zhang et al. (2024c); Garg et al. (2025)*] ] [Syntax-Semantic Space, fill=cyan!30 [Hosking and Lapata (2021)] ] [Hierarchical Syntactic Space, fill=cyan!30 [Hosking et al. (2022)] ] [Topic Attribute Space, fill=cyan!30 [Yoo and Choi (2024)] ] ] [Variational AutoEncoder (VAE), fill=green!40 [Reasoning Attribute Space, fill=green!20 [Natural Language, fill=green!10 [Zhang et al. (2024d, 2025b)] ] [Others (e.g., Vision), fill=green!10 [Sun et al. (2018); Bonnet and Macfarlane (2024); van Krieken et al. (2025a)] ] ] [Semantic Attribute Space, fill=green!20 [Carvalho et al. (2023); Zhang et al. (2024b); Zhang et al. (2025a)] ] [Syntax-Semantic Attribute Space, fill=green!20 [Zhang et al. (2019); Bao et al. (2019); Huang et al. (2021); Chen et al. (2019); Zhang et al. (2024e)] ] [Syntax Attribute Space, fill=green!20 [Shi et al. (2021); Huang and Chang (2021); Mercatali and Freitas (2021); Felhi et al. (2022)] ] [Style-transfer Attribute Space, fill=green!20 [Single Attribute, fill=green!10 [Fang et al. (2019); John et al. (2019); Hu et al. (2017); Xu et al. (2021); Mai et al. (2020);
Huang et al. (2020); Li et al. (2022); Yoshioka et al. (2022); Liu et al. (2020, 2023);
Duan et al. (2020); Shen et al. (2020); Zhang et al. (2024a)*] ] [Multiple Attributes, fill=green!10 [Gu et al. (2023, 2022); Hu and Li (2021); Vasilakes et al. (2022)] ] ] ] ]
Table 2: Taxonomy for latent semantic geometry via autoencoder architecture. For VAE and VQVAE, * indicates configurations that are plug-and-play inside LMs, rather than considering the LM as a single decoder.
@@ 4.1.1 Transformer Language VAEs
(2) KV Addition Injection. In addition to memory setup, $z$ can also be directly incorporated into the key and value matrices of the self-attention mechanism, as demonstrated in Zhang et al. (2024e).
@@
(3) Query Addition Injection. $z$ can also be directly added into the Query matrices of the self-attention mechanism for effectively search the pretrained, memorised knowledge from Key and Value by incorporating prior information Morris et al. (2025). For example, Zhang et al. (2024e) inject syntax information into Query for effective searching the semantic information from Key and Value.
@@
Latent injection setup & Math formula Summarisation
KV Addition: softmax(Q [z + K]Td) [z + V] Simple and global interventions in the sequence of attention of LM decoder, which can mitigate the KL vanishing without additional parameters Zhang et al. (2024e).
@@ 4.1.2 Latent Space Geometry
(3) Syntax-Semantic Attribute Space. Several studies have explored methods for learning latent spaces in which syntactic structure and lexical semantics are disentangled. For example, Zhang et al. (2019) used two separate encoders to independently encode semantic and syntactic information. In contrast, Bao et al. (2019) utilised a single encoder within a VAE framework to disentangle semantic and syntactic representations through multi-task optimisation. Additionally, Zhang et al. (2024e) employed a graph-based encoder alongside a semantic encoder to explicitly capture syntax and semantics, mitigating the information bottleneck from a single encoder.
(4) Semantic Attribute Space. Defining generative factors in terms of semantic roles has recently been explored as a promising direction for structuring latent spaces. Carvalho et al. (2023) defined the generative factors of natural language definition, named Definition Semantic Roles (DSR) Silva et al. (2018), to assist the Definition Modelling tasks. Recent studies Zhang et al. (2024b, 2025a) formalise the latent space geometry from the perspective of Argument Structure Theory Jackendoff (1992); Levin (1993); Rappaport Hovav and Levin (2008). In this framework, each semantic role–word content combination is represented as a convex cone within the latent space, enabling localised semantic control and enhancing interpretability.
(5) Reasoning Attribute Space. More recent studies have also explored the encoding of reasoning patterns within the latent space of VAEs for Natural Language Inference (NLI) tasks Zhang et al. (2024d, 2025b). They seek to structure the latent space in a way that captures inferential relationships, such as mathematical reasoning rules, thereby enabling the model to represent not only the meanings of individual sentences but also the logical relations that hold between them.
@@ 4.2.1 Vector Quantized VAEs
(1) Topic Attribute Space. Topic-VQVAE Yoo and Choi (2024) focuses on document-level generation by learning topic-aware latent codes. It leverages vector quantization to discover discrete topic representations that guide the generation process, enabling more coherent and thematically consistent document outputs.
@@
(4) Semantic Space. To investigate the geometrical properties of discrete latent spaces, a recent study by Zhang et al. (2024c) integrates the VQ-VAE framework into the T5 encoder-decoder architecture to learn a discrete semantic space. The authors further propose techniques such as traversal and interpolation to explore and manipulate representations within this discrete space, targetting fine-grained semantic control.
@@ 5.2 Generation Evaluation
The interpolation mechanism operates by selecting two input sentences, $x_{1}$ and $x_{2}$, and encoding them into their corresponding latent representations $z_{1}$ and $z_{2}$. A linear interpolation is then performed along the latent space trajectory defined by $z_{t}=z_{1}\cdot(1-t)+z_{2}\cdot t$, where $t$ varies from 0 to 1 in increments of 0.1. Consequently, nine intermediate sentences are generated along the interpolation path. Prior work has shown that improved latent space geometry leads to smoother and more semantically coherent interpolations Zhang et al. (2024c).
@@
In VAE, for example, given two sentences that share the concept animal—animals require food for survival and animals require warmth for survival—their vector addition in the latent space may yield a new sentence such as animals produce milk, which also preserves the concept animal. This suggests the existence of a “convex cone” corresponding to the semantic concept animal within the latent space Zhang et al. (2024c, 2025a). Moreover, using arithmetic to steer the task-specific generative factor, such as sentiment, has been widely deployed in language VAE domain Shen et al. (2020); Mercatali and Freitas (2021). Furthermore, controlling (Large) LMs generation through the arithmetic of steering vectors to internal representations has been widely explored Turner et al. (2023), such as Concept Activation Addition (CAA) Rimsky et al. (2024), Representation Perturbation (RePe) Zou et al. (2023), Top PC Im and Li (2025), and Inference-Time Intervention (ITI) Li et al. (2023). In this context, the discrete concept vectors learned by SAEs can be interpreted as steering vectors, enabling targeted modulation of language model outputs Chalnev et al. (2024), such as SPARE Zhao et al. (2025) and SAE-SSV He et al. (2025).
@@ 6 Conclusion and Open Questions
Emerging diffusion-based LMs, such as Neuro-symbolic diffusion Nie et al. (2025); van Krieken et al. (2025b), offer new opportunities to combine the generative strengths of large-scale models with symbolic-level controllability. However, the design of semantically interpretable latent priors within these frameworks is still in its early stages. Incorporating the compositional semantic representation can potentially improve the sampling efficiency during the denoising process. Developing structured latent representations that align with formal semantics could enable more transparent reasoning and controllable generation in future generative architectures.
@@ Appendix A Additional Examples
Figure 4: Generation evaluation for VAE or VQVAE latent space. More results can be found in Zhang et al. (2024b) for VAE and Zhang et al. (2024c) for VQVAE.